OpenStudy (anonymous):

Write an equation and a solution for the problem. A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound?

OpenStudy (mertsj):

let b = pounds of bran flakes let m = pounds of oatmeal b+m=30 3.6m+4.8b=4(30)